OpenStudy (anonymous):

How do I simplify the radical expression of the square root of 1 over 4??

OpenStudy (anonymous):

\[\sqrt{1/4} ..?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{a/b} =\sqrt{a}/\sqrt{b}\]

OpenStudy (anonymous):

\[\sqrt{1}/\sqrt{4}=1/2\]

OpenStudy (anonymous):

oh, okay, I had thought that it would have been 1 over 16.

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

my pleasure:D

OpenStudy (anonymous):

it becomes 1 over 16 when you need to square (1/4)

OpenStudy (anonymous):

oh, okay